RC Meeting, Monday, October 25 , 2021
RC Meeting Information
RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920
- Mostly steady production on day/swing/owl. Beam was off 09:45-13:30 on Sunday for a magnet issue that only impacted Hall C.
- The accelerator still has an unresolved issue regarding A beam going thru the C slit and causing BLM trips in Arc 7. Hall C beam is sometimes reduced so Hall A can go to higher currents.
- Regarding the Secretary of Energy tour, VIP schedules can change unexpectedly, but the last info we were given was that the accelerator will be off 09:30 Monday morning. Tours wrap up at 13:30.
- Magnet teams will try to isolate the short in the Arc 7 magnet.
- Pass change for Hall A is today/Monday, starting after the accelerator is locked up again.
Hall C Status
- We had good production in the last 24 hours, mostly at 40-42muA. Hall C beam current was reduced to 30muA for a couple hours when Hall A was increased from 3 to 6 muA.
- Number of good pions was 5139 as of 08:00 Monday. See Petr's summary at the bottom of https://logbooks.jlab.org/entry/3931960 .
Current Run Plan
- Day shift only: Try to take 1 hour of Dummy target each day. (a 100% live hour)
- Tentatively scheduled: a BCM calibration on Wednesday mid-afternoon.
- Continue data taking until early Nov
- Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
- Q3 insulating vacuum spikes generating alarms. For example, https://logbooks.jlab.org/entry/3930175 .
- One HB LHe incident of note: https://logbooks.jlab.org/entry/3929152 . (possible clue to the origin?)
- Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
- Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
- HCCAM02 not functioning (lift required to replace).